OpenStudy (anonymous):

According to a recent edition of the Guinness Book of World Records the highest rotary speed ever attained was 2010 m/s. This speed occurred at one end of a rod 15.3 cm long that rotated about the other end. What was the frequency of rotation of the rod in revolutions per minute?

OpenStudy (irishboy123):

just play with these but keep the units right...."rpm" \[v = \omega r\] \[w = \frac{2 \pi}{T} = 2 \pi f\]